# From Efficiency Gains to Rebound Effects: The Problem of Jevons’ Paradox in AI’s Polarized Environmental Debate

### v1

@@ 4. Discussion
As mentioned in Section 3, one under-explored dimension in these debates involves the indirect or rebound effects that AI can generate. While much attention has been given to AI improving productivity and resource efficiency, these gains can result in higher overall consumption due to these effects. This paradox can manifest in several ways: For instance, an AI-driven logistics system might reduce delivery times and fuel usage per vehicle, yet simultaneously encourage more frequent online orders, thus elevating total miles driven; AI-driven targeted advertising can help us more easily find the products that we need based on our clicks and searches, but also push us to make more superfluous purchases. Such systemic shifts in behavior challenge linear expectations that efficiency improvements alone will drive down emissions. Instead, they underscore the need for a detailed, interdisciplinary approach that links AI deployment to broader assessments of environmental, social, and economic feedback loops.
@@
Yet, as mentioned in Section 3.4, for any of these steps toward meaningful change to materialize, the industry must adopt a far more transparent stance on all the environmental impacts of AI systems and take accountability for the far-reaching impacts of the technologies that it develops and deploys. At present, the scant public information on the carbon footprint of large-scale AI models is frequently derived from academic estimates or limited corporate disclosures rather than comprehensive, standardized reporting. If AI companies truly want to position their technologies as part of the climate solution, they must be forthcoming with granular data on their energy sources, resource consumption, hardware lifecycles, and the end-of-life management of electronic components. In other words, the foundation of any net-positive AI contribution to the environment is a baseline of reliable, detailed data, which has yet to be made widely available. Without this level of transparency, policymaking bodies, researchers, and the public at large are left with partial insights at best, which undermines the capacity to assess AI’s environmental impacts effectively or to design incentives that reward lower-impact AI development. Ultimately, the AI field is responsible for knowing the impacts of its own products, and it cannot do so without better data. This information is crucial for performing accurate lifecycle assessments that capture both direct and indirect consequences. Absent such data, the conversation around AI’s climate benefits risks devolving into corporate branding exercises rather than a genuine reckoning with environmental and social responsibilities.
@@ 5. Conclusion
This paper argues that the AI field needs to adopt a more detailed and nuanced approach to framing, articulating, and addressing AI’s environmental impacts in order to avoid unhelpful polarization. This requires including AI’s direct impacts—mineral supply chain studies, carbon emissions of training large-scale models, water consumption, and e-waste from hardware—as well as mapping the ways AI innovations reshape economic structures and societal practices that, in turn, drive increased resource usage. Such a comprehensive perspective will empower researchers, policymakers, and industry stakeholders to devise strategies that prevent “tech-solutionism” from overshadowing the urgent need for systemic change. Greater transparency in reporting energy usage, more robust lifecycle assessment tools, and meaningful industry-wide enforceable standards are examples that would foster much-needed progress.

### v2

@@ 4. Discussion
As mentioned in Section 3, one under-explored dimension in these debates involves the indirect or rebound effects that AI can generate. While much attention has been given to AI improving productivity and resource efficiency, these gains can result in higher overall consumption due to effects such as Jevons Paradox. This paradox can manifest in several ways: for instance, an AI-driven logistics system might reduce delivery times and fuel usage per vehicle, yet simultaneously encourage more frequent online orders, thus elevating total miles driven; AI-driven targeted advertising can help us more easily find the products that we need based on our clicks and searches, but also push us to make more superfluous purchases. Such systemic shifts in behavior challenge linear expectations that efficiency improvements alone will drive down emissions. Instead, they underscore the need for a detailed, interdisciplinary approach that links AI deployment to broader assessments of environmental, social, and economic feedback loops.
Applying Jevons’ Paradox to technologies like AI has notable conceptual and empirical limitations due to the complexity inherent in technological development and diffusion (Sorrell and Dimitropoulos, 2008; Saunders, 2009). While Jevons’ Paradox anticipates increases in resource consumption following efficiency improvements, its explanatory power diminishes within the broader systemic contexts of widespread AI diffusion. Sorrell (2008) critiques Jevons’ paradox for oversimplifying complex causal relationships among efficiency gains, economic expansion, and individual behavioral responses. This is particularly relevant given AI systems have many interdependencies across diverse economic sectors. Addressing these multifaceted rebound effects requires shifting from simplistic causal assumptions toward more integrated sociotechnical analyses that can include a range of approaches, from detailed empirical studies of technology adoption, user behavior, policy factors, and evolving market dynamics. That would also suggest that truly effective policy responses for AI sustainability would need to combine technological efficiency with institutional reforms and incentive structures specifically designed to decouple model development from unsustainable resource consumption. This approach would not only recognize the conditional validity of Jevons’ Paradox but actively embed sustainability within AI development.
@@
Yet, as mentioned in Section 3.4, for any of these steps toward meaningful change to materialize, the industry must adopt a far more transparent stance on all the environmental impacts of AI systems and take accountability for the far-reaching impacts of the technologies that it develops and deploys. At present, the scant public information on the carbon footprint of large-scale AI models is frequently derived from academic estimates or limited corporate disclosures rather than comprehensive, standardized reporting. If AI companies truly want to position their technologies as part of the climate solution, they must be forthcoming with granular data on their energy sources, resource consumption, hardware lifecycles, and the end-of-life management of electronic components. In other words, the foundation of any net-positive AI contribution to the environment is a baseline of reliable, detailed data, which has yet to be made widely available. Without this level of transparency, policymaking bodies, researchers, and the public at large are left with partial insights at best, which undermines the capacity to assess AI’s environmental impacts effectively, to design incentives that reward lower-impact AI development, and for individuals to make informed choices with respect to their use of these technologies. Ultimately, the AI field is responsible for knowing the impacts of its own products, and it cannot do so without better data. This information is crucial for performing accurate lifecycle assessments that capture both direct and indirect consequences. Absent such data, the conversation around AI’s climate benefits risks devolving into corporate branding exercises rather than a genuine reckoning with environmental and social responsibilities.
@@ 5. Conclusion
This paper argues that the AI field needs to adopt a more detailed and nuanced approach to framing, articulating, and addressing AI’s environmental impacts in order to avoid unhelpful polarization. This requires including AI’s direct impacts—mineral supply chain studies, carbon emissions of training large-scale models, energy and water consumption, and e-waste from hardware—as well as mapping the ways AI innovations reshape economic structures and societal practices that, in turn, drive increased resource usage. Such a comprehensive perspective will empower researchers, policymakers, and industry stakeholders to devise strategies that prevent “tech-solutionism” from overshadowing the urgent need for systemic change. Greater transparency in reporting energy usage, more robust lifecycle assessment tools, and meaningful industry-wide enforceable standards are examples that would foster much-needed progress.